Shadee Corp. expects to sell 600 sun visors in May and 330 in June. Each visor sells for $19. Shadee’s beginning and ending finished goods inventories for May are 70 and 45 units, respectively. Ending finished goods inventory for June will be 65 units.

It expects the following unit sales for the third quarter:      

   
July 560
August 470
September 460
 


Sixty percent of Shadee’s sales are cash. Of the credit sales, 52 percent is collected in the month of the sale, 38 percent is collected during the following month, and 10 percent is never collected.

Required:

Calculate Shadee’s total cash receipts for August and September. (Do not round your intermediate calculations. Round your answers to the nearest whole dollar.)
Answer #1

SOLUTION

Cash Receipts for August-

Cash Sales Collected in August - 470 * $19 = $8,930 * 60% = $5,358

Credit sales collected in August - 470 * $19 = $8,930 * 40% = $3,572 * 52% = 1,857.44

July sales collected in August - 560 *$19 = $10,640 *40% = $4,256 * 38% = $1,617.28

Cash Collection in August = $5,358 + 1,857.44 + $1,617.28 = $8,833

Cash Receipts for September-

Cash Sales Collected in September - 460 * $19 = $8,740 * 60% = $5,244

Credit sales collected in September - 460 * $19 = $8,740 * 40% = $3,496 * 52% = $1,817.92

August sales collected in September - 470 * $19 = $8,930 * 40% = $3,572 * 38% = $1,357.36

Cash Collection in September = $5,244 + $1,817.92 + $1,357.36 = $8,419